For the following scores: 5, 3, 6, 2, 2, 2, 8 Find the deviation for each score, and check that the deviations add up to zero. X Deviation 5 3 6 2 2 2 8 sum
Answer #1

Average = Sum/n = (5 + 3 + 6 + 2 + 2 + 2 + 8)/7 = 28/7 = 4

Hence,

X Deviation
5 5 - 4 = 1
3 3 - 4 = -1
6 6 - 4 = 2
2 2 - 4 = -2
2 2 - 4 = -2
2 2 - 4 = -2
8 8 - 4 = 4
Sum 1 - 1 + 2 - 2 - 2 - 2 + 4 = 0